The sum of two numbers is 45. if three time the larger number is increased by four times the smaller number the sum is 155. Find the numbers.
Answer by checkley79(3341)

Let x be the smaller number.
Let y be the larger number.
x+y=45
3y+4x=155
x=45-y
3y+4(45-y)=155
3y+180-4y=155
-y=155-180
-y=-25
y=25 ans.
x+25=45
x=45-25
x=20 ans.
PROOF:
3*25+4*20=155
75+80=155
155=155